OpenStudy (loser66):

How to prove there is no branch of logarithm on C \{0} Help me, please

OpenStudy (loser66):

@ikram002p @dan815

OpenStudy (anonymous):

if there were such a branch, it would be the anti derivative of \(\frac{1}{z}\)

OpenStudy (loser66):

You mean log z?

OpenStudy (anonymous):

yes

OpenStudy (loser66):

but if we take z is in real and negative, then log z is not continuous on the region, right?

OpenStudy (loser66):

*undefined

OpenStudy (anonymous):

you have a theorem that says something like if \(f\) is analytic on a simply connected region then \[\int_{\gamma}f(z)dz=0\]

OpenStudy (loser66):

No! I didn't study integral yet.

OpenStudy (anonymous):

oh then this is not going to work sorry

OpenStudy (anonymous):

dang

OpenStudy (empty):

Wait, there isn't a branch of log(z) on C\{0} ?

OpenStudy (anonymous):

here, this is a more long winded explanation, see 4

OpenStudy (loser66):

where is 4 to see? :)

OpenStudy (loser66):

oh, problem 21, right?

OpenStudy (loser66):

Question: What is the link between G and G'?

OpenStudy (anonymous):

it's in there somewhere as are the details,

OpenStudy (loser66):

Why is G' subset of G? I am sorry for being dummy. I don't get it

OpenStudy (anonymous):

oh now i have to explain it too ? i thought cheating was good enough' ok lets see, seems that \(G'\) .. actually that is a real good question

OpenStudy (loser66):

cheating is good enough but understanding the solution is better. :)

OpenStudy (anonymous):

OH i didn't read carefully because \(z<0\) made no sense, but upon more careful reading it says \(z\in \mathbb{R}\)so it is negative real numbers

OpenStudy (loser66):

I think I got why G' is subset of G. Thanks for the link.

OpenStudy (loser66):

This is my interpretation, { negative number } > {0} , hence C \ {negative number}< C -{0}

OpenStudy (anonymous):

oh it is a subset because it is the negative real numbers are a subset of the complex numbers without 0

OpenStudy (anonymous):

the gimmick seems to be taking the limit as \(z=\to -1\) along the real line from the left and from the right, see that you get different answers

OpenStudy (loser66):

I got the leftover, just wonder how G' is subset of G. Again , thanks a ton.

OpenStudy (anonymous):

yw i think the real idea, (if i remember correctly) is that in order to have a branch of the log, you have to cut the complex plane, not just delete one point

OpenStudy (anonymous):

btw short proof is that it would be the anti derivative of \(\frac{1}{z}\) making \[\int_{|z|=1 }\frac{dz}{z}=0\] but it isn't

OpenStudy (loser66):

You mean the restrict region must be a line, not a point?